;
DateSerial No.Notes
2017-09-29noneHeimann mirror lens cleaned using long implement from underneath.
2016-01-21noneNo signal from the Heimann on the test flights, though calibrator signal was present and moving as it should. Traced to a D-25 plug on the rear core console, front face, window side towards the bottom, which was cut just behind the bulkhead plug at last C-Check. This was labelled SATCOM C on both plug and socket, though was in fact the Heimann signal (1-signal (red); 2-signal-return (blu); 14-screen (blk)). Both now relabelled. Crimp pins connected and inserted in new plug on bulkhead. Plug used was MH connectors one that I had in a drawer (RS 719-5798 (shell) & RS 719-5824 (crimp)).
2015-01-07876Instrument range changed from [-20 to +80] to [-50 to +50]
2013-04-01876Instrument range changed from [-50 to +50] to [-20 to +80]
2013-03-05877NPL Calibrated both cameras and the blackbody PRT
2013-03-05876NPL Calibrated both cameras and the blackbody PRT
2012-01-09noneHeimann Calibrator target found unserviceable, Target temperature displays as flashing 49.9C, obviously over-range. Target cables belled though ok, problem found to be actually at PRT where terminals had come unsoldered. Target PRT replaced with spare, having similar resistance at room temperature (within 0.5 ohms). 'Calibrations' now possible, but absolute accuracy unknown.
2009-10-27H686BAcamera and control unit
2009-10-07H686BAresults very similar to 2005 cal in B125 lab using water bath
2009-10-05H686BAfor cal
2009-04-06H686BAWhole rack refitted after rebuild.
2009-03-11H686BAWhole rack removed.
2008-01-30H686BBLoaded calibration programs and PID control settings into Eurotherm. Tested on aircraft
2008-01-16H686BAHeimann and calibration control unit removed and tested in lab. Worked fine. Aircraft cables belled through. Worked fine. Replaced on Aircraft and tested. Worked fine. Fault attributed to bad/loose/dirty connection.
2008-01-16H686BANULL
2008-01-14H686BAFor troubleshooting.
2007-11-21H686BAAt QA Lab, Exeter
2007-11-16H686BANULL
2007-10-05H686BBEntire rack removed.
2007-01-24H686BBTarget has no temperature control. Inspected and reseated pressure box connectors - appeared okay. No 115V light on Control Unit. Checked all fuses - okay. Remade all CU connectors. Switched on and it is now working!
2006-11-06H686BBNULL